# On diffusion-based generative models and their error bounds: The log-concave case with full convergence estimates

### v4

@@ 3.2 General Case
In this section, we derive the full non-asymptotic estimates in Wasserstein distance of order two between the target data distribution $\pi_{\mathsf{D}}$ and the generative distribution of the diffusion model under the assumptions stated below. As explained in Section 2 (see also Appendix A), it could be necessary in the general setting to restrict $t\in[\epsilon,T]$ for $\epsilon\in(0,1)$ in (6). Therefore, we truncate the integration in the backward diffusion at $t_{K}=T-\epsilon$ and run the process $(Y_{t})_{t\in[0,T-\epsilon]}$, $(Y_{t}^{\text{aux}})_{t\in[0,T-\epsilon]}$ and $(Y^{\text{EM}}_{k})_{k\in\{0,\dots,K\}}$ up to time $T-\epsilon$.
@@ 3.2.1 Assumptions for the General Case
In the motivating example in Section 3.1, we have chosen the SGLD algorithm to solve the optimisation problem (6). Other algorithms, such as ADAM (Kingma and Ba, 2015), TheoPouLa (Lim and Sabanis, 2024) and stochastic gradient descent (Jentzen et al., 2021), can be chosen as long as they satisfy the following assumption. Fix $\epsilon>0$.
@@
Let $\theta^{*}$ be a minimiser of (6) and let $\hat{\theta}$ be the (random) estimator of $\theta^{*}$ obtained through some approximation procedure such that $\mathbb{E}[|\hat{\theta}|^{4}]<\infty$. There exists $\widetilde{\varepsilon}_{\text{AL}}>0$ such that
@@
Remark 3.4.
As a consequence of Assumption 2 and the preservation of strong log-concavity under convolution, see, e.g., (Saumard and Wellner, 2014, Proposition 3.7), there exists $L_{\text{MO}}:[0,T]\rightarrow(0,\infty]$ such that for all $t\in[0,T]$ and $x,\bar{x}\in\mathbb{R}^{\text{M}}$, we have
@@
The function $L_{\text{MO}}(t)$ in (19) has a lower bound for all $t\in[0,T]$, which we denote by $\widehat{L}_{\text{MO}}$. Moreover, Assumption 2 with the estimate (19) implies that the processes in (3) and in (4) have a unique strong solution, see, e.g., (Krylov, 1991, Theorem 1).
Next, we consider the following assumption on the approximating function $s$ and its gradient.
Assumption 3.
@@
Moreover, there exists $\mathsf{K}_{4}>0$ such that, for all $x,\bar{x}\in\mathbb{R}^{M}$ and for any $k=1,\dots M$,
@@
Remark 3.5.
Let $\mathsf{K}_{\text{Total}}:=\mathsf{K}_{1}+\mathsf{K}_{2}+\mathsf{K}_{3}+|s(0,0,0)|>0$. Using Assumption 3, we have
@@
We postpone the proof of Remark 3.5 to Appendix C.1.
Remark 3.6.
Assumption 3 imposes Lipschitz continuity on a family of approximating functions $s(\cdot,\cdot,\cdot)$ with respect to the input variables, $t$ and $x$, as well as the parameters $\theta$. It is well-known that the continuity properties of neural networks with respect to $t$, and $x$ are largely determined by the activation function at the last layer. For instance, neural networks with activation functions such as hyberbolic tangent and sigmoid functions at the last layer satisfy the Lipschitz continuity with respect to $t$ and $x$ (Virmaux and Scaman, 2018; Fazlyab et al., 2019).
For the following assumption on the score approximation, we let $\hat{\theta}$ be as in Assumption 1 and we let $(Y_{t}^{\text{aux}})_{t\in[0,T]}$ be the auxiliary process defined in (9).
@@
We highlight that the expectation in Assumption 4 is taken over the auxiliary process (9). This density is known since the approximating function $s$ and the estimator $\hat{\theta}$ are known. To the best of authors’ knowledge, this is a novelty with respect to previous works (Bortoli, 2022; Chen et al., 2023d; Lee et al., 2022, 2023; Chen et al., 2023a; Conforti et al., 2023; Benton et al., 2024) which consider the unknown density of the forward process (or its numerical discretization).
Remark 3.8.
The motivating example in Section 3.1 is a special case of the general setting. Indeed, Assumption 1 is satisfied due to Lemma B.2 with $\hat{\theta}=\theta_{n}^{\lambda}$ being the $n$th-iterate of the SGLD algorithm (14) and $\theta^{*}:=\mu$. Assumption 2 is satisfied by the score function (12). Assumption 3 is satisfied by the approximating function (13) with $d=M$, $\alpha=1$, $D_{1}(\theta,\bar{\theta}):=|\bar{\theta}|$, $D_{2}(t,\bar{t}):=e^{-t}$, $D_{3}(t,\bar{t}):=1$, for any $\theta$, $\bar{\theta}\in\mathbb{R}^{d}$, $t$, $\bar{t}\in[0,T]$ and $\mathsf{K}_{1}=\mathsf{K}_{2}=\mathsf{K}_{3}=\mathsf{K}_{4}=1$. Furthermore, both functions satisfy Assumption 4 with $d=M$. Indeed, by Assumption 1 with $\theta^{*}=\mu$, we have
Remark 3.9.
@@
We conduct a numerical experiment to show the convergence of diffusion-based generative models under Assumption 1, 2, 3 and 4. We consider the case where $X_{0}\sim\pi_{\mathsf{D}}=\mathscr{N}(\mu,I_{d})$ with $\mu=(-1.2347,-0.89244)$ and $d=2$. We use SGLD algorithm (15) to solve the optimization problem (7), with $\kappa(t)=\sigma_{t}^{2}$, $\epsilon=0$, $T=2$, $\lambda=5\times 10^{-5}$, $\beta=10^{12}$, and $n=4\times 10^{4}$. At each iteration, $128$ mini-batch samples are used to estimate the stochastic gradient. Then, we generate samples using the Euler-Maruyama approximation (10) with $\gamma=10^{-3}$ and $s$ given in (13).
At each iteration $n$, we evaluate the quality of $100,000$ generated samples using the Wasserstein distance of order two. In addition, we compute the $L^{2}$ error between the score function and the approximated function $s$ with $\theta_{n}^{\lambda}$, using the auxiliary process as in Assumption 4. Figure 1 (a) demonstrates the error in Assumption 4 vanishes as the generative model converges, where the degree of convergence is measured by $W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})$. This empirical observation justifies Assumption 4. Moreover, we explore the relationship between the quality of generated samples and the error using (7) via denoising score matching in Figure 1 (b).
@@
(b) Error using $U(\theta)$ in (7).
Figure 1: The quality of generated samples with respect to (a) the error with Assumption 4 and (b) the error obtained through denoising score matching using $U(\theta)$ in (7).
@@ 3.2.2 Full Estimates for the General Case
The main result under the general setting is stated as follows. An overview of the proof can be found in Appendix C.2.
Theorem 3.10.
Let Assumptions 1, 2, 3 and 4 hold. Then, there exist constants $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}>0$ such that for any $T>0$ and $\gamma,\epsilon\in(0,1)$,
$\begin{split}W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})&\leq C_{1}\sqrt{\epsilon}+C_{2}e^{-2\widehat{L}_{\text{MO}}(T-\epsilon)-\epsilon}+C_{3}(T,\epsilon)\sqrt{\varepsilon_{\text{SN}}}+C_{4}(T,\epsilon)\gamma^{\alpha},\end{split}$ (21)
where $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}$ are given explicitly in Table 3 (Appendix D). In addition, the result in (21) implies that for any $\delta>0$, if we choose $0\leq\epsilon<\epsilon_{\delta}$, $T>T_{\delta}$, $0<\varepsilon_{\text{SN}}<\varepsilon_{\text{SN},\delta}$ and $0<\gamma<\gamma_{\delta}$ with $\epsilon_{\delta}$, $T_{\delta}$, $\varepsilon_{\text{SN},\delta}$, and $\gamma_{\delta}$ given in Table 3, then
$W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})<\leavevmode\nobreak\ \delta.$
Remark 3.11.
The error bounds in (21) are not as good as the ones provided in Theorem 3.1 due to the general form of the approximating function $s$. We emphasize that the optimal rate of convergence of order $\alpha\in\left[\frac{1}{2},1\right]$ for the Euler or Milstein scheme of SDEs with constant diffusion coefficients is achieved using the Lipschitz continuity on the derivative of $s$ in Assumption 3. In the explicit expression of $C_{4}$ in Table 3, the dependence of the dimension is $O(M)$ due to numerical techniques from Kumar and Sabanis (2019) used in the proof of Theorem 3.10 to achieve the optimal rate of convergence.
@@ 4 Related Work and Comparison
We describe assumptions and results of various existing works in our notation and framework to facilitate the comparisons with our results which are provided in 2-Wasserstein distance. Beyond being theoretical relevant, the choice of the use of this metric is motivated by its applications in generative modeling. A popular performance metric currently used to examine the quality of the images produced by generative models is the Fréchet Inception Distance (FID) which was originally introduced in (Heusel et al., 2017). This metric measures the Fréchet distance between the distribution of generated samples and the distribution of real samples, assuming Gaussian distributions, which is equivalent to the Wasserstein distance of order two. Our results provided under this metric are therefore also relevant for practical applications.
@@ 4.1 Score Approximation Assumptions
Most of the recent analysis of score-based generative models, see, e.g., (Chen et al., 2023d; Lee et al., 2022, 2023; Chen et al., 2023a; Conforti et al., 2023; Benton et al., 2024), have considered assumptions (in $L^{2}$) on the absolute error of the following type, i.e. for any $k\in\{0,\dots,N-1\}$, there exists $\underline{\varepsilon}>0$
@@
where the expectation is taken with respect to the unknown $\{p_{t}\}_{t\in[0,T]}$ and $\hat{\theta}$ is a deterministic quantity. In (Chen et al., 2023a; Conforti et al., 2023; Benton et al., 2024), the assumption (22) is written in integral (averaged) form. In (Lee et al., 2023), the bound in (22) is not uniform over $t$, i.e. $\underline{\varepsilon}_{t}:=\frac{\underline{\varepsilon}}{\sigma_{t}^{4}}$ and this allows the score function to grow as $\frac{1}{\sigma_{t}^{2}}$ as $t\rightarrow 0$. The authors in (Conforti et al., 2023) use the relative $L^{2}$-score approximation error such as
@@
where the expectation is with respect to the unknown density of the law of $X_{t}$ against the Gaussian distribution $\pi_{\infty}(x)$, i.e. $\tilde{p}_{t}(x):=p_{t}(x)/\pi_{\infty}(x)$, and $\tilde{s}(t,\theta,x)$ in (23) is the approximating function for the score function of $\tilde{p}_{t}$. A pointwise assumption in (Bortoli, 2022), given by
@@
for $C\geq 0$, is considered under the manifold (compact) setting. The assumption (24) takes into account the explosive behaviour of the score function as $t\rightarrow 0$. In an attempt to obtain a weaker control than (24), the assumption (Bortoli, 2022, A5) is used, namely
@@
We note that, unlike (22) and (23), the expectation in (25) is taken with respect to the algorithm $Y_{k}^{\text{EI}}$ given by the exponential integrator (EI) discretization scheme [^1] This analysis can be extended to the Euler-Maruyama numerical scheme in $Y_{k}^{\text{EM}}$ (10)., which has known density. However, the bounds of (Bortoli, 2022, Theorem H.1) in Wasserstein distance of order one derived under this assumption scale exponentially in problem parameters such as the diameter of the manifold $\mathfrak{M}$ and the inverse of the early stopping time parameter $\epsilon$, i.e. $\exp(O(\text{diam}(\mathfrak{M})^{2}/\epsilon))$. Furthermore, an assumption similar to (25) is considered in a very recent and concurrent result [^2] The concurrent paper (Gao et al., 2023) appeared few days earlier when the first draft of this work was made available online. (Gao et al., 2023):
@@
We emphasize that the existing results in the literature do not take the expectation with respect to the stochastic optimizer $\hat{\theta}$. This, together with the use of our novel auxiliary process (9) that uses only known information, allows us to deduce state-of-the-art bounds in the Wasserstein distance of order two in the following sense. The bounds scale polynomially in the data dimension $M$, i.e. $O(\sqrt{M})$ in the motivating example (Theorem 3.1) and $O(M)$ in the general case (Theorem 3.10) while achieving the optimal rate of convergence $\gamma$ in both cases.

### v5

@@ 3.2 General Case
In this section, we derive the full non-asymptotic estimates in Wasserstein distance of order two between the target data distribution $\pi_{\mathsf{D}}$ and the generative distribution of the diffusion model under the assumptions stated below. As explained in Section 2 (see also Appendix A), it could be necessary in the general setting to restrict $t\in[\epsilon,T]$ for $\epsilon\in(0,1)$ in 6. Therefore, we truncate the integration in the backward diffusion at $T-\epsilon$ and run the process $(Y_{t})_{t\in[0,T-\epsilon]}$.
@@ 3.2.1 Assumptions for the General Case
In the motivating example in Section 3.1, we have chosen the SGLD algorithm to solve the optimisation problem 6. Other algorithms, such as ADAM (Kingma & Ba, 2015), TheoPouLa (Lim & Sabanis, 2024) and stochastic gradient descent (Jentzen et al., 2021), can be chosen as long as they satisfy the following assumption. Fix $\epsilon>0$.
@@
Let $\theta^{*}$ be a minimiser [^1] The score-matching optimization problem 6 is not necessarily (strongly) convex. of 6 and let $\hat{\theta}$ be the (random) estimator of $\theta^{*}$ obtained through some approximation procedure such that $\mathbb{E}[|\hat{\theta}|^{4}]<\infty$. There exists $\widetilde{\varepsilon}_{\text{AL}}>0$ such that
@@
Remark 4.
As a consequence of Assumption 2 and the preservation of strong log-concavity under convolution, see, e.g., Saumard & Wellner (2014, Proposition 3.7), there exists $L_{\text{MO}}:[0,T]\rightarrow(0,\infty]$ such that for all $t\in[0,T]$ and $x,\bar{x}\in\mathbb{R}^{\text{M}}$, we have
@@
The function $L_{\text{MO}}(t)$ in 19 has a lower bound for all $t\in[0,T]$, which we denote by $\widehat{L}_{\text{MO}}$. Moreover, Assumption 2 with the estimate 19 implies that the processes in 3 and in 4 have a unique strong solution, see, e.g., Krylov (1991, Theorem 1).
Next, we consider the following assumption on the approximating function $s$ which is used in Remark 12.
Assumption 3.a.
@@
By adding a further condition on the gradient of $s$ in Assumption 3.a we are able to achieve the optimal rate of convergence in Theorem 10 below.
Assumption 3.b.
Let $s$ be as in Assumption 3.a and there exists $\mathsf{K}_{4}>0$ such that, for all $x,\bar{x}\in\mathbb{R}^{M}$ and for any $k=1,\dots M$,
@@
Remark 5.
Let $\mathsf{K}_{\text{Total}}:=\mathsf{K}_{1}+\mathsf{K}_{2}+\mathsf{K}_{3}+|s(0,0,0)|>0$. Using Assumption 3.b, we have
@@
We postpone the proof of Remark 5 to Appendix D.1.
Remark 6.
Assumption 3.a and 3.b impose Lipschitz continuity on a family of approximating functions $s(\cdot,\cdot,\cdot)$ with respect to the input variables, $t$ and $x$, as well as the parameters $\theta$. It is well-known that the continuity properties of neural networks with respect to $t$, and $x$ are largely determined by the activation function at the last layer. For instance, neural networks with activation functions such as hyberbolic tangent and sigmoid functions at the last layer satisfy the Lipschitz continuity with respect to $t$ and $x$ (Virmaux & Scaman, 2018; Fazlyab et al., 2019).
For the following assumption on the score approximation, we let $\hat{\theta}$ be as in Assumption 1 and we let $(Y_{t}^{\text{aux}})_{t\in[0,T]}$ be the auxiliary process defined in 9.
@@
We highlight that the expectation in Assumption 4 is taken over the auxiliary process 9. This density is known since the approximating function $s$ and the estimator $\hat{\theta}$ are known. To the best of authors’ knowledge, this is a novelty with respect to previous works (Bortoli, 2022; Chen et al., 2023d; Lee et al., 2022; 2023; Chen et al., 2023a; Conforti et al., 2025; Benton et al., 2024) which consider the unknown density of the forward process (or its numerical discretization).
Remark 8.
Assumption 4 is satisfied, along with Assumption 3.b, for data distributions satisfying Assumption 2, beyond the multivariate Gaussian case discussed in the motivating example. Indeed,
$\begin{split}&\mathbb{E}\int_{0}^{T-\epsilon}|\nabla\log p_{T-r}(Y_{r}^{\text{aux}})-s(T-r,\hat{\theta},Y_{r}^{\text{aux}})|^{2}\ \text{d}r\\ &\leq 2\ \mathbb{E}\int_{0}^{T-\epsilon}|\nabla\log p_{T-r}(Y_{r}^{\text{aux}})-s(T-r,\theta^{*},Y_{r}^{\text{aux}})|^{2}\ \text{d}r\\ &\quad+2\ \mathbb{E}\int_{0}^{T-\epsilon}|s(T-r,\theta^{*},Y_{r}^{\text{aux}})-s(T-r,\hat{\theta},Y_{r}^{\text{aux}})|^{2}\ \text{d}r.\end{split}$ (21)
If $\theta^{*}$ is such that $s(t,\theta^{*},x)=\nabla\log p_{t}(x)$ as in the motivating example in Section 3.1, then the first term on the right-hand side of 21 vanishes. Otherwise, we expect that the first term on the right-hand side above to be small. Indeed, by the definition of strong log-concavity, see e.g. Saumard & Wellner (2014, Definition 2.9), we have
$\begin{split}\nabla\log p_{t}(x)&=\nabla\log(g(x))+\nabla\log(\phi_{t}(x)),\end{split}$ (22)
where $g$ is some log-concave function and $\phi_{t}$ is a multivariate Gaussian density. If $g$ is a multivariate logistic distribution [^2] The multivariate logistic distribution is an example of elliptical distribution widely used in portfolio risk management (Xiao & Valdez, 2015; Owen & Rabinovitch, 1983). (Malik & Abraham, 1973), then its score function is given by
$\begin{split}\frac{\partial}{\partial x_{k}}\log(g(x))&=-1-(M+1)\frac{-\exp(-x_{k})}{1+\sum_{k=1}^{M}\exp(-x_{k})},\qquad\quad-\infty<x_{k}<\infty,\qquad k=1,\dots,M,\end{split}$ (23)
while the Hessian of $\log(g(x))$ is bounded (see Appendix B for more details). Therefore, the score function of the multivariate logistic distribution 23 is Lipschitz, and, as a consequence, $\nabla\log p_{t}(x)$ in 22 is still Lipschitz. Thus, we expect to have a good control on the first term on the right-hand side of 21 since the function $s$ satisfying Assumption 3.b approximates a Lipschitz score function. In general, there exists a function
$c(t,x):=\nabla\log p_{t}(x)-s(t,\theta^{*},x),\qquad t>0,\ x\in\mathbb{R}^{M},$ (24)
and therefore one has to define a log-concave function $g$ such that the first term on the right-hand side of 21 is small. Clearly, this is a problem specific challenge and it may not always have a good solution. The second term on the right-hand side of 21 is controlled by Assumption 3.b and Assumption 1 (see Appendix B for more details).
Remark 9.
@@
We conduct a numerical experiment to show the convergence of diffusion-based generative models under Assumption 1, 2, 3.b and 4. We consider the case where $X_{0}\sim\pi_{\mathsf{D}}=\mathcal{N}(\mu,I_{d})$ with $\mu=(-1.2347,-0.89244)$ and $d=2$. We use SGLD algorithm 15 to solve the optimization problem 7, with $\kappa(t)=\sigma_{t}^{2}$, $\epsilon=0$, $T=2$, $\lambda=5\times 10^{-5}$, $\beta=10^{12}$, and $n=4\times 10^{4}$. At each iteration, $128$ mini-batch samples are used to estimate the stochastic gradient. Then, we generate samples using the Euler-Maruyama approximation 10 with $\gamma=10^{-3}$ and $s$ given in 13. At each iteration $n$, we evaluate the quality of $100,000$ generated samples using the Wasserstein distance of order two. In addition, we compute the $L^{2}$ error between the score function and the approximated function $s$ with $\theta_{n}^{\lambda}$, using the auxiliary process as in Assumption 4. Figure 1 (a) demonstrates the error in Assumption 4 vanishes as the generative model converges, where the degree of convergence is measured by $W_{2}(\mathcal{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})$. This empirical observation justifies Assumption 4. Moreover, we explore the relationship between the quality of generated samples and the error using 7 via denoising score matching in Figure 1 (b).
@@
(b) Error using $U(\theta)$ in 7.
Figure 1: The quality of generated samples with respect to (a) the error with Assumption 4 and (b) the error obtained through denoising score matching using $U(\theta)$ in 7.
@@ 3.2.2 Full Estimates for the General Case
The main result under the general setting is stated as follows. An overview of the proof can be found in Appendix D.2.
Theorem 10.
Let Assumptions 1, 2, 3.b and 4 hold. Then, there exist constants $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}>0$ such that for any $T>0$ and $\gamma,\epsilon\in(0,1)$,
$\begin{split}W_{2}(\mathcal{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})&\leq C_{1}\sqrt{\epsilon}+C_{2}e^{-2\widehat{L}_{\text{MO}}(T-\epsilon)-\epsilon}+C_{3}(T,\epsilon)\sqrt{\varepsilon_{\text{SN}}}+C_{4}(T,\epsilon)\gamma^{\alpha},\end{split}$ (25)
where $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}$ are given explicitly in Table 4 (Appendix E). In addition, the result in 25 implies that for any $\delta>0$, if we choose $0\leq\epsilon<\epsilon_{\delta}$, $T>T_{\delta}$, $0<\varepsilon_{\text{SN}}<\varepsilon_{\text{SN},\delta}$ and $0<\gamma<\gamma_{\delta}$ with $\epsilon_{\delta}$, $T_{\delta}$, $\varepsilon_{\text{SN},\delta}$, and $\gamma_{\delta}$ given in Table 4, then
$W_{2}(\mathcal{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})<~{}\delta.$
Remark 11.
The error bounds in 25 are not as good as the ones provided in Theorem 1 due to the general form of the approximating function $s$. We emphasize that the optimal rate of convergence of order $\alpha\in\left[\frac{1}{2},1\right]$ for the Euler or Milstein scheme of SDEs with constant diffusion coefficients is achieved using the Lipschitz continuity on the derivative of $s$ in Assumption 3.b. In the explicit expression of $C_{4}$ in Table 4, the dependence of the dimension is $O(M)$ due to numerical techniques from Kumar & Sabanis (2019) used in the proof of Theorem 10 to achieve the optimal rate of convergence.
Remark 12.
If we replace Assumption 3.b with Assumption 3.a in Theorem 10, then the bound 25 becomes
$\begin{split}W_{2}(\mathcal{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})&\leq C_{1}\sqrt{\epsilon}+C_{2}e^{-2\widehat{L}_{\text{MO}}(T-\epsilon)-\epsilon}+C_{3}(T,\epsilon)\sqrt{\varepsilon_{\text{SN}}}+\widetilde{C}_{4}(T,\epsilon)\gamma^{1/2},\end{split}$
where $C_{1}$, $C_{2}$, $C_{3}$ are the same as in Theorem 10 and $\widetilde{C}_{4}(T,\epsilon)$ contains a better dependence on the dimension, namely $O(\sqrt{M})$, than the one achieved in Theorem 10, and it is explicitly provided in Table 4 (Appendix E). Although this relaxation achieves the same dependence on the data dimension as in the motivating example in Theorem 1, it leads to a worse rate of convergence of order $1/2$.
@@ 4 Related Work and Comparison
We describe assumptions and results of various existing works in our notation and framework to facilitate the comparisons with our results which are provided in 2-Wasserstein distance. Beyond being theoretical relevant, the choice of the use of this metric is motivated by its applications in generative modeling. A popular performance metric currently used to examine the quality of the images produced by generative models is the Fréchet Inception Distance (FID) which was originally introduced in Heusel et al. (2017). This metric measures the Fréchet distance between the distribution of generated samples and the distribution of real samples, assuming Gaussian distributions, which is equivalent to the Wasserstein distance of order two. Our results provided under this metric are therefore also relevant for practical applications.
@@ 4.1 Score Approximation Assumptions
The early work of De Bortoli et al. (2021) requires an $L^{\infty}$-bound on the score approximation, which is in contrast with the $L^{2}$ nature of the score-matching optimization problem 6. Most of the recent analysis of score-based generative models, see, e.g., Chen et al. (2023d); Lee et al. (2022; 2023); Chen et al. (2023a); Conforti et al. (2025); Benton et al. (2024), have considered assumptions (in $L^{2}$) on the absolute error of the following type, i.e. for any $k\in\{0,\dots,N-1\}$, there exists $\underline{\varepsilon}>0$
@@
where the expectation is taken with respect to the unknown $\{p_{t}\}_{t\in[0,T]}$ and $\hat{\theta}$ is a deterministic quantity. In Chen et al. (2023a); Conforti et al. (2025); Benton et al. (2024), the assumption 26 is written in integral (averaged) form. In Lee et al. (2023), the bound in 26 is not uniform over $t$, i.e. $\underline{\varepsilon}_{t}:=\frac{\underline{\varepsilon}}{\sigma_{t}^{4}}$ and this allows the score function to grow as $\frac{1}{\sigma_{t}^{2}}$ as $t\rightarrow 0$. The authors in Conforti et al. (2025) use the relative $L^{2}$-score approximation error such as
@@
where the expectation is with respect to the unknown density of the law of $X_{t}$ against the Gaussian distribution $\pi_{\infty}(x)$, i.e. $\tilde{p}_{t}(x):=p_{t}(x)/\pi_{\infty}(x)$, and $\tilde{s}(t,\theta,x)$ in 27 is the approximating function for the score function of $\tilde{p}_{t}$. A pointwise assumption in Bortoli (2022), given by
@@
for $C\geq 0$, is considered under the manifold (compact) setting. The assumption 28 takes into account the explosive behaviour of the score function as $t\rightarrow 0$. In an attempt to obtain a weaker control than 28, the assumption Bortoli (2022, A5) is used, namely
@@
We note that, unlike 26 and 27, the expectation in 29 is taken with respect to the algorithm $Y_{k}^{\text{EI}}$ given by the exponential integrator (EI) discretization scheme [^3] This analysis can be extended to the Euler-Maruyama numerical scheme in $Y_{k}^{\text{EM}}$ 10., which has known density. However, the bounds of Bortoli (2022, Theorem H.1) in Wasserstein distance of order one derived under this assumption scale exponentially in problem parameters such as the diameter of the manifold $\mathfrak{M}$ and the inverse of the early stopping time parameter $\epsilon$, i.e. $\exp(O(\text{diam}(\mathfrak{M})^{2}/\epsilon))$. Furthermore, an assumption similar to 29 is considered in a very recent and concurrent result [^4] The concurrent paper Gao et al. (2023) appeared few days earlier when the first draft of this work was made available online. Gao et al. (2023):
@@
We emphasize that the existing results in the literature do not take the expectation with respect to the stochastic optimizer $\hat{\theta}$. This, together with the use of our novel auxiliary process 9 that uses only known information, allows us to deduce state-of-the-art bounds in the Wasserstein distance of order two in the following sense. The bounds scale polynomially in the data dimension $M$, i.e. $O(\sqrt{M})$ as shown in Theorem 1 and Remark 12, and achieve the optimal rate of convergence: of order one in Theorem 1 and of order $\alpha\in\left[\frac{1}{2},1\right]$ in Theorem 10.